OpenStudy (angel_kitty12):

I forgot how to solve 256 -1/4 Can someone help please?

Vocaloid (vocaloid):

is the -1/4 an exponent?

OpenStudy (angel_kitty12):

Sorry it's 256^ -1/4 The fraction is the exponent

Vocaloid (vocaloid):

256^(-1/4) = 1/(256^(1/4)) = 1/4

OpenStudy (anonymous):

256 ^ -1/4 = 256 ^ -.25 so basically divide 256 by 4 and make it negative

OpenStudy (angel_kitty12):

But where did the 256 go?

OpenStudy (angel_kitty12):

Wait never mind I understand

OpenStudy (anonymous):

You got it? Good! What did you get?

OpenStudy (angel_kitty12):

Thank you very much :)